# RoSTE: An Efficient Quantization-Aware Supervised Fine-Tuning Approach for Large Language Models

### v2

@@ Abstract
Supervised fine-tuning is a standard method for adapting pre-trained large language models (LLMs) to downstream tasks. Quantization has been recently studied as a post-training technique for efficient LLM deployment. To obtain quantized fine-tuned LLMs, conventional pipelines would first fine-tune the pre-trained models, followed by post-training quantization. This often yields suboptimal performance as it fails to leverage the synergy between fine-tuning and quantization. To effectively realize low-bit quantization of weights, activations, and KV caches in LLMs, we propose an algorithm named Rotated Straight-Through-Estimator (RoSTE), which combines quantization-aware supervised fine-tuning (QA-SFT) with an adaptive rotation strategy that identifies an effective rotation configuration to reduce activation outliers. We provide theoretical insights on RoSTE by analyzing its prediction error when applied to an overparameterized least square quantized training problem. Our findings reveal that the prediction error is directly proportional to the quantization error of the converged weights, which can be effectively managed through an optimized rotation configuration. Experiments on Pythia, Qwen and Llama models of different sizes demonstrate the effectiveness of RoSTE. Compared to existing post-SFT quantization baselines, our method consistently achieves superior performances across various tasks and different LLM architectures. Our code is available at https://github.com/Hong-Lab-UMN-ECE/RoSTE.
@@ 1 Introduction
While fine-tuned models excel in domain-specific tasks, their substantial computational and storage demands present challenges for efficient deployment, particularly in resource-constrained environments (Xu et al., 2024a). To address these limitations, various model compression techniques have been developed, including quantization (Lin et al., 2023; Frantar et al., 2022), pruning (Ma et al., 2023; Sun et al., 2023), distillation (Xu et al., 2024b), and low-rank approximation (Wang et al., 2024a; Yuan et al., 2023). Among these, quantization is particularly effective for compressing LLMs, as it significantly reduces memory consumption, inference latency, and power usage. Additionally, its compatibility with specialized hardware accelerators enhances its practical deployment across diverse devices. Quantization techniques generally fall into two categories: post-training quantization (PTQ) and quantization-aware training (QAT). PTQ is well-suited for quick deployment with minimal resources but often sacrifices accuracy in low-bit settings. In contrast, QAT achieves effective compression with minimal performance loss but requires retraining the entire LLM on a large corpus, incurring substantial computational costs.
@@
This work presents one of the first studies on quantization-aware fine-tuning (QA-SFT) to obtain effective fine-tuned and quantized LLM as a single training phase. To maximize the hardware capability of modern GPUs, we concentrate on designs utilizing the 4-bit quantization of weights, activations, and KV cache in LLMs. However, low-bit quantization presents a major challenge: performance degradation due to the weight and activation outliers, which expand the quantization range and increase the quantization error. Notice that even the high-performance data-free QAT methods (Liu et al., 2023) fail with 4-bit activation quantization.
The first key aspect of our work is to leverage rotation-based quantization in QA-SFT. Our work is inspired by recent findings on rotation-based PTQ methods (Ashkboos et al., 2024b; Liu et al., 2024), which demonstrate that applying offline and online rotations to linear projection layers and KV caches in LLMs effectively mitigates weight and activation outliers in post-trained models. However, when directly applied to QA-SFT, these PTQ methods fail to prevent outliers from re-emerging within layers during fine-tuning, resulting in performance degradation. To address this, we propose a joint training method combining the adaptive selection of rotation matrices and QA-SFT. The second key aspect of our work is to utilize a bilevel optimization formulation that simultaneously tackles QA-SFT and selects the rotation matrices based on the weights.
@@
  • We provide a theoretical analysis of the benefits of rotation-enabled quantization in QA-SFT by examining the prediction error resulted from the QAT stage of RoSTE. This analysis directly motivates the use of quantization error based surrogate loss and justifies the adoption of the low-complexity Walsh-Hadamard rotation.
We conduct experiments on fine-tuning Pythia, Qwen and Llama models, demonstrating the effectiveness of RoSTE. An accuracy-vs-training-time plot in Figure 1 illustrates the practical benefit of RoSTE for obtaining a fine-tuned quantized LLM. Overall, we believe this is the first work that develops an efficient quantization algorithm for the SFT process, with theoretical justification and practical effectiveness.
@@ 2 Preliminary
This section introduces ingredients that are essential to the proposed RoSTE algorithm through overviewing two major approaches for achieving efficient quantized LLMs.
@@ 2.1 Post-Training Quantization (PTQ)
with $c\in(0,1]$ a clipping factor used to scale down the representation range so as to mitigate the impact of outlier values. To take advantage of the representation range in tensor with value distribution skewed away from 0, we can adopt asymmetric quantization by choosing
@@ 2.2 Quantization-Aware Training (QAT)
The main objective of quantization-aware training (QAT) is to directly optimize a quantized neural network using gradient-based methods. From an optimization perspective, this is challenging as the quantization operator is not differentiable. To this regard, we focus on a popular remedy that has been studied before the era of LLMs is the straight-through estimator (STE) (Courbariaux et al., 2015; Bai et al., 2018) which approximates the Jacobian of quantizer by the identity matrix. During the backward calculation, the derivative of quantizer $Q$ in the chain rule is replaced by
@@ 2.3 Supervised Fine-Tuning (SFT)
Although it is a natural idea to apply QAT on fine-tuning tasks, existing works such as (Dettmers et al., 2023; Xu et al., 2023) (also see (Lee et al., 2024; Bondarenko et al., 2024) for similar ideas applied to training LLMs) only considered quantization aware adaptation utilizing an additional LoRA architecture. Moreover, they focused on direct quantization without incoherence processing whose performance can be sensitive to outliers. This motivates us to consider integrating QAT with incoherence processing for SFT.
@@ 3 Proposed Algorithm: RoSTE
This section presents the Rotated Straight-Through-Estimator (RoSTE) algorithm through jointly optimizing the rotated quantizer and model parameters. To fix the idea, we parameterize the quantized LLM by the weight matrices $\{{\bf W}_{i}\}_{i=0}^{\ell-1}$ and rotation matrices $\{\mathbf{R}_{i}\}_{i=0}^{\ell-1}$. Consider the abstraction of an LLM with $\ell$ layers/modules as $\mathbf{m}_{Q}:\overline{\cal X}\to\mathbb{R}^{|{\cal T}|}$, where $\overline{\cal X}$ is the set of sequences with variable context length, ${\cal T}$ is the set of vocabulary, we denote
@@
${\cal R}^{k}=\operatorname*{\arg\!\min}_{\mathbf{R}_{0},...,\mathbf{R}_{\ell-1}\in\{{\bf H},{\bf I}\}}{\cal E}({\cal W}^{kT},\{\mathbf{R}_{i}\}_{i=0}^{\ell-1}),$ (9)
where ${\bf H}$ is a random Walsh-Hadamard matrix generated according to (19).
@@
11:        Output: Quantized fine-tuned $\mathbf{m}_{Q}(~{}\cdot~{};{\cal W}^{KT},{\cal R}^{K-1})$. However, tackling (8) can be challenging even with approaches such as alternating optimization. This is because, upon fixing the weight matrices, minimizing the objective function w.r.t. the rotation matrices $\{\mathbf{R}_{i}\}_{i=0}^{\ell-1}$ involves a non-tractable manifold optimization while the objective function is non-differentiable due to quantization. Meanwhile, when the rotation matrices are fixed, the minimization problem w.r.t. the weight matrices $\{{\bf W}_{i}\}_{i=0}^{\ell-1}$ is similar to standard QAT; see (Liu et al., 2023). The above obstacle motivated us to consider an alternative formulation (albeit somewhat heuristic) that simplifies the search for $\{\mathbf{R}_{i}\}_{i=0}^{\ell-1}$ adapted to the weight matrices. This formulation explicitly separates the process of (quantized) model training and the rotation matrix optimization, and leverages a simpler loss function for the rotation matrix design. More specifically, we consider:
@@
for $\mathbf{X}_{i,j}$ representing the input activation of layer $i$ on the $j$-th calibration data sample, e.g., by drawing a subset of size $n$ from ${\cal D}$ the fine-tuning dataset. Notice that now the optimal lower level variable aims at assisting the upper level weights so that an STE gradient approximation on ${\cal L}_{\rm SFT}$ w.r.t. $\{{\bf W}_{i}\}_{i=0}^{\ell-1}$ has a smaller bias. However, it remains challenging for us to access the optimal rotation matrices for every iteration of the upper level minimization as solving the lower level problem can still be computationally expensive. In this regard, we propose a lazy lower level approximation where the rotation matrices are updated after $T$ iterations of optimizing the weight matrices.
@@
The RoSTE algorithm is now summarized in Algorithm 3 and Fig. 2. The algorithm is akin to alternating optimization and consists of two parts. The first part (cf. line 79) pertains to the QAT stage with SFT objective for selecting the weight matrices $\{{\bf W}_{i}\}_{i=0}^{\ell-1}$ under the rotation matrices. Notice that STE gradient can be computed efficiently when $\{\mathbf{R}_{i}\}$ are chosen as the Walsh-Hadamard matrices; also see (15). To our best knowledge, the subroutine used in this stage is the first direct application of STE to SFT training as well as its adaptation to the possibly varying rotation matrices. The second part (cf. line 5) pertains to the selection of rotation matrices in the lower level optimization which is a non-smooth problem on the manifold. Compared to (8), its objective function can be easily computed as the latter only represents the quantization error. Furthermore, as we will show in Sec. 4, the random Walsh-Hadamard matrix ${\bf H}$ yields an approximate-but-universal solution to minimize ${\cal E}(\cdot)$. As such, we propose to approximate the subproblem by limiting the search space to $\mathbf{R}_{i}\in\{{\bf H},{\bf I}\}$ using a random Hadamard matrix ${\bf H}$ (Tseng et al., 2024) and perform a (low-complexity) combinatorial search to obtain an approximate lower level solution that decides if the rotation matrix should be applied on each layer. Details of this heuristic implementation can be found in Appendix D.
@@ 4 Theoretical Insights of RoSTE
Our setup follows from the literature on analyzing the convergence of SGD for neural networks under the interpolation regime (Ma et al., 2018; Vaswani et al., 2019). To describe it, let us fix the rotation matrices $\{\mathbf{R}_{i}\}_{i=0}^{\ell-1}$ and consider the QAT stage (cf. line 79) in the RoSTE algorithm. Instead of analyzing ${\cal L}_{\rm SFT}(\mathbf{m}_{Q}(\cdot))$ directly, we consider the quadratic loss function as a simplified objective to draw insights for RoSTE. Moreover, the training dataset consists of samples $(\mathbf{x}_{\xi},{\bf y_{\xi}})$ with a single output token in $\mathbb{R}$ such that $\mathbf{y}_{\xi}\in{\cal Y}\equiv\mathbb{R}$. For any $\mathbf{m}:\overline{\cal X}\to\mathbb{R}^{|{\cal T}|}$, we now consider the squared prediction error:
@@
We further assume that the composite map $o(\mathbf{m}_{Q}(\cdot))$ is a linear quantized model given by
@@
The above assumption requires that the quantized-rotated features $(\mathbf{x}_{\xi},\mathbf{y}_{\xi})$ are interpolatable by a full-precision model $\mathbf{w}^{\star}_{\mathbf{R}}$. This assumption is closely related to the standard interpolation assumption that appeared in the literature on training with over-parameterized models (Ma et al., 2018; Vaswani et al., 2019). It is worth noticing that Assumption 4.2 does not require the interpolator $\mathbf{w}_{\mathbf{R}}^{\star}$ to be in the quantized model parameter space (14).
@@
for any $t\geq 0$, where $\mu=\frac{\lambda_{\rm min}^{2}}{12\rho}$ and ${\bf e}(\mathbf{x})=Q_{w}(\mathbf{x})-\mathbf{x}$.
@@
Suppose that the QAT stage of RoSTE is run with $T\gg 1$ inner-loop iterations. Applying the theorem shows that given $\overline{\mathbf{R}}$, the resultant prediction error of the intermediate model $\mathbf{w}^{T}$ will be bounded by ${\cal O}(\sum_{s=0}^{T}\left(1-\mu\right)^{T-s}\mathbb{E}\left[{\|Q_{w}(\overline{\mathbf{R}}\mathbf{w}^{s})-\overline{\mathbf{R}}\mathbf{w}^{s}\|_{\mathbf{G}}^{2}}\right])$, i.e., a weighted sum of the weight quantization errors during the QAT process. Due to the exponential weighting $\left(1-\mu\right)^{T-s}$, the prediction error is dominated by the weight quantization error of recent iterates. Crucially, the above analysis shows that the rotation matrices play a pivoting role in the performance of QAT. This inspires us to apply ${\cal E}(\cdot)$ in (12) to guide us in the selection for optimal rotation matrices.
@@
The analysis in (22) enables a novel interpretation of the bit-widths in $Q_{x}$ and $Q_{w}$ during STE training. On one hand, it is beneficial to increase the bit-width of activation quantization $Q_{x}$ until Assumption 4.2 is satisfied, and further increasing its bit-width would not improve the prediction performance as the bound (18) only depends on weight quantization error. On the other hand, increasing the bit-width of weight quantization always reduces the prediction error as seen in (18), (21). It is also interesting to see that despite adopting low-bit activation quantizers, increasing the dimension $d$ may still allow us to satisfy the interpolation condition Assumption 4.2. This is due to the intuition that kernelized high dimensional features are more likely to be separable (Liang & Rakhlin, 2020). In other words, a neural network with high-dimensional hidden representations can tolerate low-bit quantized activations because the information about $\mathbf{x}_{\xi}$ retains in the high-dimensional discrete vector $Q_{x}(\mathbf{R}\mathbf{x}_{\xi})$.

### v3

@@ Abstract
Supervised fine-tuning is a standard method for adapting pre-trained large language models (LLMs) to downstream tasks. Quantization has been recently studied as a post-training technique for efficient LLM deployment. To obtain quantized fine-tuned LLMs, conventional pipelines would first fine-tune the pre-trained models, followed by post-training quantization. This often yields suboptimal performance as it fails to leverage the synergy between fine-tuning and quantization. To effectively realize low-bit quantization of weights, activations and KV caches in LLMs, we propose an algorithm named Rotated Straight-Through-Estimator (RoSTE), which combines quantization-aware supervised fine-tuning (QA-SFT) with an adaptive rotation strategy that identifies an effective rotation configuration to reduce activation outliers. We provide theoretical insights on RoSTE by analyzing its prediction error when applied to an overparameterized least square quantized training problem. Our findings reveal that the prediction error is directly proportional to the quantization error of the converged weights, which can be effectively managed through an optimized rotation configuration. Experiments on Pythia, Qwen and Llama models of different sizes demonstrate the effectiveness of RoSTE. Compared to existing post-SFT quantization baselines, our method consistently achieves superior performances across various tasks and different LLM architectures. Our code is available at https://github.com/OptimAI-Lab/RoSTE.
Machine Learning, ICML
@@ 1 Introduction
While fine-tuned models excel in domain-specific tasks, their substantial computational and storage demands present challenges for efficient deployment, particularly in resource-constrained environments (Xu et al., 2024a). To address these limitations, various model compression techniques have been developed, including quantization (Lin et al., 2023; Frantar et al., 2022), pruning (Ma et al., 2023; Sun et al., 2023), distillation (Xu et al., 2024b), and low-rank approximation (Wang et al., 2024a; Yuan et al., 2023). Among these, quantization is particularly effective for compressing LLMs, as it significantly reduces memory consumption, inference latency and power usage. Additionally, its compatibility with specialized hardware accelerators enhances its practical deployment across diverse devices. Quantization techniques generally fall into two categories: post-training quantization (PTQ) and quantization-aware training (QAT). PTQ is well-suited for quick deployment with minimal resources but often sacrifices accuracy in low-bit settings. In contrast, QAT achieves effective compression with minimal performance loss but requires retraining the entire LLM on a large corpus, incurring substantial computational costs.
@@
This work presents one of the first studies on quantization-aware supervised fine-tuning (QA-SFT) to obtain effective fine-tuned and quantized LLM through a single training phase. To maximize the hardware capability of modern GPUs, we concentrate on designs utilizing the 4-bit quantization of weights, activations, and KV cache in LLMs. However, low-bit quantization presents a major challenge on models with weight and activation outliers: they expand the quantization range and increase the quantization error, degrading the quantized model accuracy. This partly explains why high-performance data-free QAT methods (Liu et al., 2023) fail at 4-bit activation quantization.
The first key aspect of our work is to leverage rotation-based quantization in QA-SFT. Our work is inspired by recent findings on rotation-based PTQ methods (Ashkboos et al., 2024b; Liu et al., 2024), which demonstrate that applying offline and online rotations to linear projection layers and KV caches in LLMs effectively mitigates weight and activation outliers in post-trained models. However, it is not clear whether one-shot PTQ appraoches should be performed before or after fine-tuning. To address this, we propose a joint training method combining an adaptive selection of rotation matrices and QA-SFT. The second key aspect of our work is to utilize a bilevel optimization formulation that simultaneously tackles QA-SFT and selects the rotation matrices based on the weights and activations.
@@
  • We provide a theoretical analysis of the benefits of rotation-enabled quantization in QA-SFT by examining the prediction error resulted from the QAT stage of RoSTE. This analysis directly motivates the use of quantization error-based surrogate loss and justifies the adoption of the low-complexity Walsh-Hadamard rotation.
We conduct experiments on fine-tuning Pythia, Qwen and Llama models, demonstrating the effectiveness of RoSTE. An accuracy-vs-training-time plot in Figure 1 illustrates that RoSTE finds quantized fine-tuned models with improved downstream tasks accuracy at the cost of a marginally longer training time. Overall, we believe this is the first work that develops an efficient quantization algorithm for the SFT process, with theoretical justification and practical effectiveness.
@@ 2 Preliminary
This section provides an overview of two major approaches for achieving efficient quantized LLMs, which are two key ingredients to the proposed RoSTE algorithm.
@@ 2.1 Post-Training Quantization (PTQ)
with $c\in(0,1]$ a constant clipping factor used to scale down the representation range so as to mitigate the impact of outlier values. To take advantage of the representation range in tensor with value distribution skewed away from 0, we can adopt asymmetric quantization by choosing
@@ 2.2 Quantization-Aware Training (QAT)
The main objective of quantization-aware training (QAT) is to directly optimize a quantized neural network using gradient-based methods. From an optimization perspective, this is challenging as the quantization operator is not differentiable. To this regard, straight-through estimator (STE) (Courbariaux et al., 2015; Bai et al., 2018) is a commonly adopted remedy which approximates the Jacobian of quantizer by the identity matrix. During the backward calculation, the derivative of quantizer $Q$ in the chain rule is replaced by
@@ 2.3 Supervised Fine-Tuning (SFT)
Although it is a natural idea to apply QAT on fine-tuning tasks, existing works such as (Dettmers et al., 2023; Xu et al., 2023) (also see (Lee et al., 2024; Bondarenko et al., 2024) for similar ideas applied to training LLMs) only considered quantization aware adaptation utilizing an additional LoRA architecture. Moreover, they focused on direct quantization without incoherence processing whose performance can be sensitive to outliers. This motivates us to consider integrating QAT with incoherence processing for SFT. In particular, we shall reveal the separate roles of weight matrices and rotation matrices from the perspective of optimization in the next section.
@@ 3 Proposed Algorithm: RoSTE
This section presents the Rotated Straight-Through-Estimator (RoSTE) algorithm through jointly optimizing the rotation matrices and model parameters. To fix the idea, we parameterize the quantized LLM by the weight matrices $\{{\bf W}_{i}\}_{i=0}^{\ell-1}$ and rotation matrices $\{\mathbf{R}_{i}\}_{i=0}^{\ell-1}$. Consider the abstraction of an LLM with $\ell$ layers/modules as $\mathbf{m}_{Q}:\overline{\cal X}\to\mathbb{R}^{|{\cal T}|}$, where $\overline{\cal X}$ is the set of sequences with variable context length, ${\cal T}$ is the set of vocabulary, we denote
@@
${\cal R}^{k}=\operatorname*{\arg\!\min}_{\mathbf{R}_{i}\in\{{\bf H},{\bf I}\}}{\cal E}({\cal W}^{kT},\{\mathbf{R}_{i}\}_{i=0}^{\ell-1}),$ (9)
where each ${\bf R}_{i}$ is chosen as either the identity matrix ${\bf I}$, i.e., no rotation, or a random Walsh-Hadamard matrix ${\bf H}$, generated according to (19).
@@
11:        Output: Quantized fine-tuned $\mathbf{m}_{Q}(~{}\cdot~{};{\cal W}^{KT},{\cal R}^{K-1})$. However, tackling (8) can be challenging even with approaches such as alternating optimization. This is because, upon fixing the weight matrices, minimizing the objective function w.r.t. the rotation matrices $\{\mathbf{R}_{i}\}_{i=0}^{\ell-1}$ involves an intractable manifold optimization while the objective function is non-differentiable due to quantization. Meanwhile, when the rotation matrices are fixed, the minimization problem w.r.t. the weight matrices $\{{\bf W}_{i}\}_{i=0}^{\ell-1}$ is similar to standard QAT; see (Liu et al., 2023). The above obstacle motivated us to consider an alternative formulation (albeit somewhat heuristic) that simplifies the search for $\{\mathbf{R}_{i}\}_{i=0}^{\ell-1}$ adapted to the weight matrices. This formulation explicitly separates the process of (quantized) model training and the rotation matrix optimization, and leverages a simpler objective function over the rotation matrices. More specifically, we consider:
@@
for $\mathbf{X}_{i,j}$ representing the input activation of layer $i$ on the $j$-th calibration data sample, e.g., by drawing a subset of size $n$ from ${\cal D}$ the fine-tuning dataset. Motivated to solving the ideal formulation (8), in our re-formulation (11), the optimal lower level variable aims at assisting the upper level weights so that an STE gradient approximation on ${\cal L}_{\rm SFT}$ w.r.t. $\{{\bf W}_{i}\}_{i=0}^{\ell-1}$ has a smaller bias. However, it remains challenging for us to access the optimal rotation matrices for every iteration of the upper level minimization as solving the lower level problem can still be computationally expensive. In this regard, we propose a lazy lower level approximation where the rotation matrices are updated after $T$ iterations of optimizing the weight matrices.
@@
The RoSTE algorithm is now summarized in Algorithm 3 and Fig. 2. The algorithm is akin to alternating optimization and consists of two parts. The first part (cf. line 79) pertains to the QAT stage with SFT objective for selecting the weight matrices $\{{\bf W}_{i}\}_{i=0}^{\ell-1}$ under the rotation matrices. Notice that the computation overhead introduced by rotations are insignificant when $\{\mathbf{R}_{i}\}$ are chosen as the Walsh-Hadamard matrices. The second part (cf. line 5) pertains to the selection of rotation matrices in the lower level optimization which is a non-smooth problem on the manifold. Compared to (8), the lower level objective function (12) can be easily computed through calculating quantized weights and a few mini-batch forward passes on sample data. Furthermore, as we will show in Sec. 4, the random Walsh-Hadamard matrix ${\bf H}$ yields an approximate-but-universal solution to minimize ${\cal E}(\cdot)$. As such, we propose to approximate the subproblem by limiting the search space to $\mathbf{R}_{i}\in\{{\bf H},{\bf I}\}$ using a random Hadamard matrix ${\bf H}$ (Tseng et al., 2024) and perform a (low-complexity) combinatorial search to obtain an approximate lower level solution that decides if the rotation matrix should be applied on each layer. Details of this heuristic implementation can be found in Appendix D.
@@ 4 Theoretical Insights of RoSTE
Our setup follows from the literature on analyzing the convergence of SGD for neural networks under the interpolation regime (Ma et al., 2018; Vaswani et al., 2019). To describe it, let us fix the rotation matrices $\{\mathbf{R}_{i}\}_{i=0}^{\ell-1}$ and consider the QAT stage (cf. line 79) in the RoSTE algorithm. Instead of analyzing ${\cal L}_{\rm SFT}(\mathbf{m}_{Q}(\cdot))$ directly, we consider the quadratic loss function as a simplified objective to draw insights for RoSTE. Moreover, the training dataset consists of samples $(\mathbf{x}_{\xi},{\bf y_{\xi}})$ with a target output token in $\mathbb{R}$ such that $\mathbf{y}_{\xi}\in{\cal Y}\equiv\mathbb{R}$.
For any $\mathbf{m}:\overline{\cal X}\to\mathbb{R}^{|{\cal T}|}$, we now consider the squared prediction error:
@@
We further assume that the composite map $o(\mathbf{m}_{Q}(\cdot))$ is a linear activation-weight quantized model given by
@@
The above assumption requires that the quantized-rotated features $(Q_{x}(\mathbf{R}\mathbf{x}_{\xi}),\mathbf{y}_{\xi})$ are interpolatable by a full-precision model $\mathbf{w}^{\star}_{\mathbf{R}}$. This assumption is closely related to the standard interpolation assumption that appeared in the literature on training over-parameterized models (Ma et al., 2018; Vaswani et al., 2019). It is worth noticing that Assumption 4.2 does not require the interpolator $\mathbf{w}_{\mathbf{R}}^{\star}$ to be in the quantized model parameter space (14).
@@
for any $t\geq 0$, where $\mu=\frac{\lambda_{\rm min}^{2}}{12\rho}$ and ${\bf e}(\mathbf{x}):=Q_{w}(\mathbf{x})-\mathbf{x}$.
@@
More specifically, suppose the QAT stage of RoSTE is run with $T\gg 1$ inner-loop iterations. Applying the theorem shows that given $\overline{\mathbf{R}}$, the resultant prediction error of model $\mathbf{w}^{T}$ will be bounded by ${\cal O}(\sum_{s=0}^{T}\left(1-\mu\right)^{T-s}\mathbb{E}\left[{\|Q_{w}(\overline{\mathbf{R}}\mathbf{w}^{s})-\overline{\mathbf{R}}\mathbf{w}^{s}\|_{\mathbf{G}}^{2}}\right])$, i.e., a weighted sum of the weight quantization errors during the QAT process. Due to the exponential weighting $\left(1-\mu\right)^{T-s}$, the prediction error is dominated by the weight quantization error of recent iterates. Crucially, the above analysis shows that the rotation matrices play a pivoting role in the performance of QAT. This inspires us to apply ${\cal E}(\cdot)$ in (12) to guide us in the selection for optimal rotation matrices, covering the weight quantization error of the rotated weight matrices $\|{\bf e}({\bf R}_{i}^{\top}{\bf W}_{i})\|^{2}$.
@@
The analysis in Theorem 4.3 and (22) enables a novel interpretation of the bit-widths in $Q_{x}$ and $Q_{w}$ during STE training. On one hand, it is beneficial to increase the bit-width of activation quantization $Q_{x}$ until Assumption 4.2 is satisfied, and further increasing its bit-width would not improve the prediction performance as the bound (18) only depends on weight quantization error. On the other hand, increasing the bit-width of weight quantization always reduces the prediction error as seen in (18), (21). It is also interesting to see that despite adopting low-bit activation quantizers, increasing the dimension $d$ may still allow us to satisfy the interpolation condition Assumption 4.2, under the intuition that kernelized high dimensional features are more likely to be separable (Liang & Rakhlin, 2020). In other words, a neural network with high-dimensional hidden representations can tolerate low-bit quantized activations because the information about $\mathbf{x}_{\xi}$ retains in the high-dimensional discrete vector $Q_{x}(\mathbf{R}\mathbf{x}_{\xi})$.